Calculate the amount in grams of potassium chloride for it to absorb heat at 450kJ whose H is at 15kJ/mol, water has zero effect in enthalpy change
1
Expert's answer
2020-04-28T12:19:28-0400

Q = n*H

n = Q/H = 450/15 = 30 mol

m = n*M = 30*(39+35.5) = 2235 g